# Conan: A Chunkwise Online Network for Zero-Shot Adaptive Voice Conversion

### v3

@@ Abstract
Zero-shot online voice conversion (VC) holds significant promise for real-time communications and entertainment. However, current VC models struggle to preserve semantic fidelity under real-time constraints, deliver natural-sounding conversions, and adapt effectively to unseen speaker characteristics. To address these challenges, we introduce Conan, a chunkwise online zero-shot voice conversion model that preserves the content of the source while matching the speaker identity of reference speech. Conan comprises three core components: 1) a Stream Content Extractor that leverages Emformer for low-latency streaming content encoding; 2) an Adaptive Style Encoder that extracts fine-grained stylistic features from reference speech for enhanced style adaptation; 3) a Causal Shuffle Vocoder that implements a fully causal HiFiGAN using a pixel-shuffle mechanism. Experimental evaluations demonstrate that Conan outperforms baseline models in subjective and objective metrics. Audio samples can be found at https://aaronz345.github.io/ConanDemo.
@@ I Introduction
Zero-shot online voice conversion (VC) aims to extract the linguistic content from source speech and, in real time, modify its speaker identity to match that of a reference speech [1]. This field has shown considerable potential for applications in real-time communication, interactive entertainment, virtual humans, and multimedia production [2]. Existing VC approaches typically employ pre-trained feature-extraction networks to extract content information from the source speech and combine it with speaker identity embeddings from the reference speech to perform chunkwise online conversion [3].
The field of VC has seen substantial advances [4]. In particular, zero‐shot VC, which aims to transfer unseen speaker identity from reference speech to generated speech, has been explored in recent years [5]. Moreover, real-time VC systems capable of operating online have been successfully deployed in low-latency scenarios using chunkwise streaming methods [6]. However, these methods still fall short of achieving the semantic fidelity, style transfer, and naturalness required to meet the growing demand for high‑quality, customizable audio experiences [7]. For chunkwise online source speech input, it is essential not only to maintain both content extraction and speech synthesis quality under strict online latency constraints but also to transfer the unseen speaker identity, which encompasses timbre and styles (like emotion and prosody) [8].
@@
  • Achieving style transfer in zero-shot scenarios is still challenging. Most VC models rely on the strong assumption that the reference voice is accessible [12] for model adaptation, which does not always hold in practice. Moreover, many VC approaches focus solely on timbre transfer by directly applying the source speech’s F0 and energy, which contradicts the goal of capturing the reference speaker identity [3]. The handful of zero-shot VC methods that do attempt to generalize to unseen speakers still struggle to balance high speaker similarity with naturalness. This shortfall arises from insufficient modeling of speaker identity, which prevents effective alignment of speaker identity with source content [6].
@@
To address these challenges above, we introduce Conan, a chunkwise online network for zero-shot adaptive voice conversion. In zero‐shot online scenarios, Conan preserves the high‐fidelity content of the source speech while matching the unseen speaker identity of reference speech, generating highly natural speech on a chunkwise basis. To achieve high-quality chunkwise streaming content extraction, we design the Stream Content Extractor, built on an Emformer architecture and trained using content representations extracted offline by HuBERT. To enable adaptive style transfer, we introduce the Adaptive Style Encoder, which employs clustering-based vector quantization to capture detailed speaker attributes from reference speech and uses an alignment attention mechanism to fuse them with content and timbre information. For high-quality natural streaming speech synthesis, we adopt a causal-convolution mel decoder and propose the Causal Shuffle Vocoder, a fully causal HiFiGAN that leverages a pixel-shuffle mechanism to eliminate checkerboard artifacts while improving computational efficiency. Experimental evaluations show that Conan outperforms baseline models in real-time content accuracy, quality, and speaker similarity. Conan can achieve a latency as low as 37 ms for the conversion on a single A100 GPU without any engineering optimizations.
@@ II-A Zero-Shot Voice Conversion
In practical scenarios, obtaining reference speech during training is often impractical. Therefore, zero-shot VC models are essential to capture and control speaker identity by disentangling content from speaker characteristics. Chou et al. employ instance normalization to separate speaker and content [14], while Wang et al. retrieve variable‐length speaker embeddings along both temporal and channel dimensions under the guidance of a pre‐trained speaker‐verification model [15]. Ebbers et al. leverage adversarial contrastive predictive coding for fully unsupervised separation of content and speaker [16]. AutoVC squeezes out speaker timbre from content embeddings using an information bottleneck [5], and VQMIVC encodes content via vector quantization, applying mutual information constraints to decorrelate speech components [4]. Some approaches obtain speaker representations through a speaker‐verification model and extract content from Automatic Speech Recognition (ASR) posteriorgrams [11]. ControlVC [2] uses pretrained encoders to get content and speaker embeddings and applies TD-PSOLA and pitch contour manipulation for time-varying speed and pitch control. NANSY trains in a fully self‐supervised manner using wav2vec features alongside Yingram [17], and LM-VC tokenizes speech into semantic tokens via HuBERT and acoustic tokens via SoundStream [18]. Although StreamVoice further advances real‐time separation and recombination of speaker and content using a language model and ASR [1], current online VC systems still exhibit considerable room for improvement in transferring unseen reference speaker identity and in accurately extracting linguistic content.
@@ II-B Online Voice Conversion
In online VC, the system must process source speech chunk by chunk and synthesize output incrementally, without access to future context. This often degrades naturalness at chunk boundaries. For streaming applications, causal processing is a critical design consideration. Hayashi et al. [19] propose a streamable version of the non‐autoregressive sequence‐to‐sequence VC model based on FastSpeech2 [20] and NAR‐S2S‐VC [21], incorporating causal convolutions and self‐attention with causal masks. FastS2S‐VC [22] learns to predict attention distributions from source speech and reference speaker indices alone, guided by a teacher model. IBF-VC [23] uses the Intermediate Bottleneck Features (IBFs) to replace Phonetic Posteriorgrams (PPGs) in the ASR encoder to capture more fine-grained prosody information, and applies non-streaming teacher guidance for the timbre leakage problem. DualVC [24] leverages dynamic masked convolution to use the within-chunk future information better. Yang et al. [6] adapt the originally offline VQMIVC model [4] for real‐time, chunk‐by‐chunk processing. ALO‐VC [25] assembles a streaming system composed of a speaker‐verification model, a streamable phonetic‐posteriorgram extractor, and an F0 extractor. StreamVC [3] demonstrates that a lightweight causal convolutional network can effectively capture soft speech‐unit representations, while StreamVoice [1] introduces autoregressive decoders to enforce temporal causality. Quamer [13] convert non-causal vocoders into causal ones by zero-padding, [13]. However, methods relying on frame-by-frame autoregressive structures degrade the quality of early segments and incur extra computational overhead, and zero-padding causal vocoders suffer from artifacts, limiting overall naturalness.
@@ II-C Style Modeling
Modeling speaking identity remains a central challenge in speech research. Prior approaches have largely relied on pre-trained models to capture only a limited set of styles, like wav2vec 2.0 [26], HuBERT [27], and WavLM [10]. Attentron [28] introduces an attention mechanism to extract speaker identity from reference samples. ZSM-SS [29] proposes a Transformer-based architecture with an external speaker encoder based on wav2vec 2.0. Daft-Exprt [30] employs a gradient reversal layer to improve reference speaker fidelity in style transfer. GenerSpeech introduces both global and local style adapters to capture diverse speaking identity [31], while Styler decomposes style into multiple levels of supervision [32]. Yang et al. [6] jointly model speaker identity and global prosody using a GST-based style token network. Mega-TTS 2 employs vector quantization for prosody encoding combined with a language model for prosody transfer [33], and NaturalSpeech 3 uses factorized vector quantization to disentangle prosodic features [34]. CosyVoice integrates x-vectors into a large language model to both disentangle and model prosody [35]. Most of these methods focus on offline speech synthesis, while our work addresses the challenge of zero-shot online voice conversion by modeling rich speaking identity and maintaining precise content alignment.
@@ III-A Overview
The architecture of Conan is shown in Figure 1 (a). Source speech is input online and fed into the system in chunks. In the Stream Content Extractor, it is extracted as content label $z_{c-i}$ for the $i$-th chunk. Since the offline HuBERT content representation format is used, each content label corresponds to a 20 ms WAV segment. The complete unseen reference speech is passed through a mel extractor to obtain reference mel $m_{rf}$, which is then input to the timbre encoder to produce a global speaker timbre embedding $z_{t}$. Additionally, $m_{rf}$ is also fed into the Adaptive Style Encoder to extract chunk-level style representations that may include emotions and prosody. By using attention to align with $z_{t}$ and $z_{c-i}$, we obtain a style representation $z_{s-i}$ that aligns with the $i$-th content chunk. Next, the causal pitch predictor estimates F0, as explicit pitch modeling greatly aids in conveying speech style [36]. The causal mel decoder then decodes the content, timbre, style, and pitch embeddings to generate the mel spectrogram $m_{gn-i}$, which is then converted to the generated speech $y_{gn-i}$ with high quality and stability via a causal shuffle vocoder. This generated speech contains the content of the source speech, the speaker identity of the reference speech. With sufficient hardware, the processing time for each chunk is shorter than the chunk’s duration, enabling chunkwise, streaming, low-latency, zero-shot online voice conversion. Note that, except for the Stream Content Extractor’s chunk-input-based design, all other components are frame-level causal models, allowing great flexibility in handling diverse chunk size settings.
@@ III-B Stream Content Extractor
$z_{c\text{-}i}=\arg\max_{1\leq j\leq J}\Bigl{[}\,\mathrm{Softmax}\bigl{(}U\,C_{i}^{(N)}\bigr{)}\Bigr{]}_{(j,t)}.$ (6)
Here, $C_{i}^{n}$ is the content feature of the $i$-th chunk at layer $n,1<n<N$, $R_{i}^{n}$ is the right‐context block, $s_{i}^{n}$ is the chunk summary computed by mean pooling, and $M_{i}^{n}$ is the memory bank from the previous chunk. Attention is computed using $(W_{q}C_{i}^{n},K_{i}^{n},V_{i}^{n})$, the output is added residually to $C_{i}^{n}$, and fed into a feedforward network $\mathrm{FFN}$. The summary $m_{i}^{\,n+1}$ is obtained by a similar Attention. Finally, $z_{c\text{-}i}$ is produced by projecting $C_{i}^{(N)}$ with $U$, applying Softmax over $J$ classes per frame $t$, and selecting the highest‐probability label for each frame. All layers remain causal by caching $M_{i}^{n}$, $K_{L,i}^{n}$, and $V_{L,i}^{n}$, enabling low‐latency, zero‐shot online content extraction.
In this attention mechanism, keys, values, and memory‐bank vectors are employed to capture contextual dependencies between data chunks. By flexibly adjusting the size of the right context, one can trade off performance against latency, setting it to zero enforces a strictly causal behavior. The sizes of the left context and memory buffer determine the overall receptive field. Leveraging our Stream Content Extractor, we achieve efficient, high‐quality streaming content extraction.
@@ III-C Adaptive Style Encoder
To comprehensively capture and transfer speaker identity from reference speech, we introduce the Adaptive Style Encoder. As shown in Figure 2, reference mel-spectrogram $m_{rf}$ is initially refined through convolutional blocks. Subsequently, a downsampling layer pools the output into chunk-level representations. In our experiments, we set the chunk size to 80 ms to incorporate styles, which may include emotion and prosody.
Next, we employ a linear projection to map the outputs $z$ into a low-dimensional latent space for efficient code index lookup [37]. The Clustering Vector Quantization (CVQ) layer [38] extracts style representations from these latent inputs, establishing an information bottleneck that effectively filters out non-style information [39]. By applying linear projection and the CVQ mechanism, we capture fine-grained speaker identity features except global timbre and content [31].
@@
[…]
@@
After generating the style embedding, we adaptively align it with the content, taking into account the influence of global timbre on style selection, by matching it to $z_{ct-i}$, the concatenation of $z_{c-i}$ and $z_{t}$ [41]. For this purpose, we introduce the Align Attention module, incorporating the Scaled Dot-Product Attention mechanism [42]. Before inputting the detailed style embedding into the attention module, we also add positional encoding embeddings. Within this module, $z_{ct-i}$ serves as the query, while the style embedding functions as both key and value. Ultimately, we get the detailed speaker identity $z_{s-i}$.
@@ III-D Causal Shuffle Vocoder
Enforcing online VC causality through frame-by-frame autoregressive methods degrades the quality of early generated segments and incurs extra computational overhead, while using zero-padding in vocoders leads to spectral artifacts. To address this, after generating the mel-spectrogram $m_{gn-i}$ using a causal mel decoder composed of causal convolutions, we build Causal Shuffle Vocoder. Based on high-quality HiFi-GAN [43], we replace standard convolutions with causal convolutions and implement upsampling via pixel shuffle, resulting in a strictly causal vocoder with high audio quality.
As illustrated in Figure 1 (c), $m_{gn-i}$ first passes through a single causal convolutional layer, producing an initial feature map of shape $(C,T_{0})$, where $C$ and $T_{0}$ represent the channel and time dimensionality, respectively. The network then comprises $N$ upsampling stages. At the $n$-th stage, a causal convolution projects the channel dimension from $C$ up to $r_{n}C$, yielding an intermediate feature map of shape $(r_{n}C,T_{n-1})$. Next, pixel shuffling rearranges these $r_{n}C$ channels into a new tensor of shape $(C,r_{n}T_{n-1})$, where $r_{n}T_{n-1}=T_{n}$, to achieve temporal upsampling. Because pixel shuffling only reorders channels without introducing information from future frames, the network’s output is frame-level causal. By choosing the upsampling factors, the final temporal resolution becomes $T_{0}\prod_{n=1}^{N}r_{n}$, recovering the original audio sampling rate.

### v4

@@ Abstract
Zero-shot online voice conversion (VC) holds significant promise for real-time communications and entertainment. However, current VC models struggle to preserve semantic fidelity under real-time constraints, deliver natural-sounding conversions, and adapt effectively to unseen speaker characteristics. To address these challenges, we introduce Conan, a chunkwise online zero-shot voice conversion model that preserves the content of the source while matching the speaker representation of reference speech. Conan comprises three core components: 1) a Stream Content Extractor that leverages Emformer for low-latency streaming content encoding; 2) an Adaptive Style Encoder that extracts fine-grained stylistic features from reference speech for enhanced style adaptation; 3) a Causal Shuffle Vocoder that implements a fully causal HiFiGAN using a pixel-shuffle mechanism. Experimental evaluations demonstrate that Conan outperforms baseline models in subjective and objective metrics. Audio samples can be found at https://aaronz345.github.io/ConanDemo.
@@ I Introduction
Voice Conversion (VC) refers to the task that aims to extract the linguistic content from the source speech and modify its speaker representation to match that of a reference speech [1]. The field of VC has seen substantial advances [2]. A particularly influential line of work is zero-shot VC, which transfers speaker representations of unseen speakers from reference speech to generated speech, without any speaker-specific fine-tuning [3]. Most current zero-shot systems typically employ pre-trained feature extraction networks to extract content information from the source speech and combine it with speaker embeddings obtained from the reference speech [4].
The growing need for real-time communication, interactive entertainment, virtual humans, and multimedia production pushes the development of online zero-shot voice conversion systems, and real-time VC systems capable of operating online have been successfully deployed in low-latency scenarios using chunkwise streaming methods [5, 6]. However, these methods still fall short of achieving the semantic fidelity, speaker similarity, and naturalness that high-quality, customizable audio experiences require [7]. For chunkwise online source speech input, it is essential not only to maintain content extraction and speech synthesis quality under strict online latency constraints, but also to transfer the unseen speaker representation, including global representation (like timbre) and detailed representation (like emotion and prosody) [8].
@@
  • Achieving style transfer in zero-shot scenarios is still challenging. Most voice conversion models are not zero-shot and many of them rely on the assumption that the reference voice is accessible for model adaptation [12], which does not always hold in practice. Moreover, many VC approaches focus solely on timbre transfer by applying the source speech’s F0 and energy; This does not capture the speaking style of the reference speaker [6]. The handful of zero-shot VC methods that do attempt to generalize to unseen speakers still struggle to balance high speaker similarity with naturalness. This shortfall arises from insufficient modeling of speaker representation, preventing effective alignment of speaker information with content [5].
@@
To address these challenges above, we introduce Conan, a Chunkwise online network for zero-shot adaptive voice conversion. In zero‐shot online scenarios, Conan preserves the high‐fidelity content of the source speech while matching the unseen speaker representation of reference speech, generating highly natural speech on a chunkwise basis. To achieve high-quality chunkwise streaming content extraction, we design the Stream Content Extractor, built on an Emformer architecture and trained using content representations extracted offline by HuBERT. To enable adaptive style transfer, we introduce the Adaptive Style Encoder, which employs clustering-based vector quantization to capture detailed speaker attributes from reference speech and uses an alignment attention mechanism to fuse them with content and timbre information. For high-quality natural streaming speech synthesis, we adopt a causal-convolutional mel decoder and propose the Causal Shuffle Vocoder, a fully causal HiFiGAN that leverages a pixel-shuffle mechanism to eliminate checkerboard artifacts of transoposed convolution. Experimental evaluations show that Conan outperforms baseline models in real-time content accuracy, quality, and speaker similarity. Conan can achieve a latency as low as 37 ms for the conversion on a single A100 GPU without any engineering optimizations.
@@ II-A Zero-Shot Voice Conversion
In practical scenarios, obtaining reference speech during training is often impractical. Therefore, zero-shot VC models are essential to capture and control speaker representation by disentangling content from speaker characteristics. Ebbers et al. [14] leverage adversarial contrastive predictive coding for fully unsupervised separation of content and speaker. AutoVC [3] squeezes out speaker timbre from content embeddings using an information bottleneck, and VQMIVC [2] encodes content via vector quantization, applying mutual information constraints to de-correlate speech components. Some approaches obtain speaker representations through a speaker‐verification model and extract content from Automatic Speech Recognition (ASR) posteriorgrams [11]. ControlVC [4] uses pretrained encoders to get content and speaker embeddings and applies TD-PSOLA and pitch contour manipulation for time-varying speed and pitch control. NANSY [15] trains in a fully self‐supervised manner using wav2vec features alongside Yingram, and LM-VC [16] tokenizes speech into semantic tokens via HuBERT and acoustic tokens via SoundStream. Although StreamVoice [1] further advances real‐time separation and recombination of speaker and content using a language model and ASR, current online VC systems still exhibit considerable room for improvement in transferring unseen reference speaker representation and in accurately extracting linguistic content.
@@ II-B Online Voice Conversion
In online VC, the system must process source speech chunk by chunk and synthesize output incrementally, without access to future context. This often degrades naturalness at chunk boundaries. For streaming applications, causal processing is a critical design consideration. Hayashi et al. [17] propose a streamable version of the non‐autoregressive sequence‐to‐sequence VC model based on FastSpeech2 [18] and NAR‐S2S‐VC [19], incorporating causal convolutions and self‐attention with causal masks. FastS2S‐VC [20] learns to predict attention distributions from source speech and reference speaker indices alone, guided by a teacher model. IBF-VC [21] uses the Intermediate Bottleneck Features (IBFs) to replace Phonetic Posteriorgrams (PPGs) in the ASR encoder to capture more fine-grained prosody information, and applies non-streaming teacher guidance for the timbre leakage problem. DualVC [22] leverages dynamic masked convolution to use the within-chunk future information. Yang et al. [5] adapt the originally offline VQMIVC model [2] for real‐time, chunk‐by‐chunk processing. ALO‐VC [23] assembles a streaming system composed of a speaker‐verification model, a streamable phonetic‐posteriorgram extractor, and an F0 extractor. StreamVC [6] demonstrates that a lightweight causal convolutional neural network can effectively capture soft speech‐unit representations, while StreamVoice [1] introduces autoregressive decoders to enforce temporal causality. Quamer et al. [13] convert non-causal vocoders into causal ones by zero-padding. However, methods relying on frame-by-frame autoregressive structures degrade the quality of early segments and incur extra computational overhead, and zero-padding causal vocoders suffer from artifacts, limiting overall naturalness.
@@ II-C Style Modeling
Modeling speaking identity remains a central challenge in speech research. Prior approaches have largely relied on pre-trained models to capture only a limited set of styles, like wav2vec 2.0 [24], HuBERT [25], and WavLM [10]. Attentron [26] introduces an attention mechanism to extract speaker representation from reference samples. ZSM-SS [27] proposes a Transformer-based architecture with an external speaker encoder based on wav2vec 2.0. Daft-Exprt [28] employs a gradient reversal layer to improve reference speaker fidelity in style transfer. GenerSpeech [29] introduces both global and local style adapters to capture diverse speaking identity, while Styler [30] decomposes style into multiple levels of supervision. Yang et al. [5] jointly model speaker identity and global prosody using a GST-based style token network. Mega-TTS 2 [31] employs vector quantization for prosody encoding combined with a language model for prosody transfer, and NaturalSpeech 3 [32] uses factorized vector quantization to disentangle prosodic features. CosyVoice [33] integrates x-vectors into a large language model to both disentangle and model prosody. Most of these methods focus on offline speech synthesis, while our work addresses the challenge of zero-shot online voice conversion by modeling rich speaking identity and maintaining precise content alignment.
@@ III-A Overview
The architecture of Conan is shown in Figure 1 (a). Source speech is input online and fed into the system in chunks. In the Stream Content Extractor, it is extracted as content label $z_{c,i}$ for the $i$-th chunk. Since the offline HuBERT content representation format is used, each content label corresponds to a 20ms audio segment. The complete unseen reference speech is passed through a mel extractor to obtain reference mel $m_{rf}$, which is then input to the convolutional timbre encoder to produce a timbre embedding $z_{t}$. Additionally, $m_{rf}$ is also fed into the Adaptive Style Encoder to extract chunk-level style representations that may include emotions and prosody. By using attention toward $z_{t}$ and $z_{c,i}$, we obtain a style representation $z_{s,i}$ that aligns with the $i$-th content chunk. Next, the causal pitch predictor estimates F0, as explicit pitch modeling greatly aids in conveying speaking styles [34]. The causal mel decoder then decodes the content, timbre, style, and pitch embeddings to generate the mel-spectrogram of the $i$-th chunk, which is then converted to the generated speech with high quality and stability via a causal shuffle vocoder. This generated speech contains the content of the source speech and the speaker representation of the reference speech. With sufficient hardware, the processing time for each chunk is shorter than the chunk’s duration, enabling chunkwise, streaming, low-latency, zero-shot online voice conversion. Note that, except for the Stream Content Extractor’s chunk-input-based design, all other components are frame-level causal models.
@@ III-B Stream Content Extractor
$z_{c,i}=\arg\max_{1\leq j\leq J}\Bigl{[}\,\mathrm{Softmax}\bigl{(}U\,C_{i}^{(N)}\bigr{)}\Bigr{]}_{(j,t)}.$ (6)
Here, $Q_{i}^{n}$, $K_{i}^{n}$, $V_{i}^{n}$ refer to the query, key, and value of the $i$-th chunk at layer $n$, $1<n<N$, $C_{i}^{n}$ is the content feature, $R_{i}^{n}$ is the right‐context block, $s_{i}^{n}$ is the chunk summary computed by mean pooling, and $M_{i}^{n}$ is the memory bank from the previous chunk. Attention is computed using $(W_{q}C_{i}^{n},K_{i}^{n},V_{i}^{n})$, the output is added residually to $C_{i}^{n}$, and fed into a feedforward network $\mathrm{FFN}$. The summary $m_{i}^{\,n+1}$ is obtained by a similar Attention, forming part of $M_{i}^{n+1}$. Finally, $z_{c,i}$ is produced by projecting $C_{i}^{(N)}$ with $U$, applying Softmax over $J$ classes per frame $t$, and selecting the highest‐probability label for each frame. All layers remain causal by caching the memory bank $M_{i}^{n}$, the key $K_{L,i}^{n}$, and the value $V_{L,i}^{n}$ from previous segments, enabling low‐latency, zero‐shot online content extraction.
In this attention mechanism, keys, values, and memory‐bank vectors are employed to capture contextual dependencies between data chunks. By flexibly adjusting the size of the right context, one can balance performance against latency; setting it to zero enforces strictly causal behavior. The sizes of the left context and memory buffer determine the overall receptive field, and through our Stream Content Extractor, we achieve efficient, high-quality streaming content extraction.
@@ III-C Adaptive Style Encoder
To comprehensively capture and transfer speaker representation from reference speech, we introduce the Adaptive Style Encoder. As shown in Figure 2, reference mel-spectrogram $m_{rf}$ is initially refined through convolutional blocks. Subsequently, a downsampling layer pools the output into chunk-level representations. In our experiments, we set the chunk size to 80 ms to incorporate styles, which may include emotion and prosody.
Next, we employ a linear projection to map the outputs $z$ into a low-dimensional latent space for efficient code index lookup [35]. The Clustering Vector Quantization (CVQ) layer [36] extracts style representations from these latent inputs, establishing an information bottleneck that effectively filters out non-style information [37]. By applying linear projection and the CVQ mechanism, we capture detailed speaker representation features except global representation and content [29].
@@
where $e$ is the selected code from the codebook, $\text{sg}(\cdot)$ is the stop-gradient operator, and $\beta$ is commitment loss hyperparameter. For code $e$ in the codebook, we directly select the closest $z^{+}$ as the positive pair and sample other farther $z^{-}$ as negative pairs. Therefore, the contrastive loss is defined as:
@@
After generating the style embedding, we adaptively align it with the content, taking into account the influence of timbre on style selection, by matching it to $z_{ct,i}$, the concatenation of $z_{c,i}$ and $z_{t}$ [39]. For this purpose, we introduce the Align Attention module, incorporating the Scaled Dot-Product Attention mechanism [40]. Before inputting the detailed style embedding into the attention module, we also add positional encoding embeddings. Within this module, $z_{ct,i}$ serves as the query, while the style embedding functions as both key and value. Finally, we get the detailed speaker representation $z_{s,i}$.
@@ III-D Causal Shuffle Vocoder
Enforcing online VC causality through frame-by-frame autoregressive methods degrades the quality of early generated segments and incurs extra computational overhead, while using zero-padding in vocoders leads to spectral artifacts. To address this, after generating the mel-spectrogram using a causal mel decoder composed of causal convolutions, we build Causal Shuffle Vocoder. Based on high-quality HiFi-GAN [41], we replace standard convolutions with causal convolutions and implement upsampling via pixel shuffle, resulting in a strictly causal vocoder with high audio quality.
As illustrated in Figure 1 (c), the reference mel-spectrogram first passes through a single causal convolutional layer, producing an initial feature map of shape $(C,T_{0})$, where $C$ and $T_{0}$ represent the channel and time dimensionality, respectively. The network then comprises $N$ upsampling stages. At the $n$-th stage, a causal convolution projects the channel dimension from $C$ up to $r_{n}C$, yielding an intermediate feature map of shape $(r_{n}C,T_{n-1})$. Next, pixel shuffling rearranges these $r_{n}C$ channels into a new tensor of shape $(C,r_{n}T_{n-1})$, where $r_{n}T_{n-1}=T_{n}$, to achieve temporal upsampling. Because pixel shuffling only reorders all channels without introducing additional information from future frames, the network’s process is frame-level causal. By carefully designing the upsampling factors, the final temporal resolution becomes $T_{0}\prod_{n=1}^{N}r_{n}$, thereby recovering the original audio sampling rate.